Copper crystallises in fcc with a unit cell length of 361 pm. What is the radius of copper atom ?

Grade:11

2 Answers

Arun
3 years ago
As given that copper crystallizes is FCC lattice (Face centred cubic). In FCC atoms are present on the corners of the cubic, unit cell as well as on the face centres of each face.
The atoms on the face diagonal will be touching each other. Let, the radius of the atom be r and edge length of the cube be a.
Face diagonal of cube =2​a
r+2r+r=2​a
r=42​a​=22​a​=22​361​ pm=128 pm
Vikas TU
3 years ago
Edge length = a = 361pm
Radius of Cu = ?
For fcc, r = a√2/4 = 361√2/4 =
= 127.63pm
or, 127pm
So, the correct answer is (a) 127pm